Naruko885:

The weight of an African elephant is 6 × 106 grams; the weight of a tiger is 3 × 105 grams. How many times heavier is the elephant than the tiger? 2 grams 20 grams 200 grams 2,000 grams

Naruko885:

@Misery Can you help?

Misery:

\(\Huge\color{Green}{I}\) \(\Huge\color{Blue}{GOT}\) \(\Huge\color{Purple}{U}\) \(\Huge\color{Red}{BRO}\)

Misery:

I hope this is of use to you @Naruko885 http://openstudy.com/updates/54ef610ae4b0819407311b14

Naruko885:

I got 300,000 but what do I do with it?

Misery:

idk

Misery:

gimmie 1 sec

Misery:

How did you get 300k?

Naruko885:

something like this 6 x 10^6 = 6,000,000 3 x 10^5 = 30,000 and 30,000 x 20 = 6,000,000 (the weight of the elephant)

Misery:

what the problem says 6 x 106 and 3 x 105.....

Misery:

you confused me

Naruko885:

look at the best response answer.

Misery:

the closest thing i come too is 200

Naruko885:

ok ill go with that.